What volume of distilled water is needed to make an 3 M solution of HCl acid using 5 moles of HCl?


1
Expert's answer
2021-04-23T08:07:22-0400

100g of conc. HCl has 37g of HCl

density is 1.2 g/ml

volume = 100g / 1.2(g/ml) = 83.3 ml = 0.0833 L

n = m/MW n = 37g /34.5(g/mol) = 1.07 moles

[HCl] = 1.07 moles / 0.0833 L = 12.8 M!

V1 x C1 = V2 x C2

12.8M x Xml = 0.025M x 100ml

X = 0.025M x 100ml / 12.8M

X = 0.2 ml of conc.HCl bring to 100 ml with ddH2O